OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

First of all, you should probably get the slope of the equation

OpenStudy (anonymous):

You need to change the equation 4x-3y=-12 into slope intercept form

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

y=4/3x-4 is your new equation

OpenStudy (anonymous):

Then do what i showed you in your last question! Let me see that you are understanding it so try this one out and i will help you if you make mistakes

OpenStudy (anonymous):

Y=4/3x-4 (-3,4) y-4=4/3x+3 (distribute) y-4=4/3x+4 add 4 y=4/3x+8

OpenStudy (anonymous):

Is this what you got?